# Rectangular recurrence relations in $\mathfrak{gl}_{n}$ and $\mathfrak{o}_{2n+1}$ invariant integrable models

### v1

@@ 3.3 Embedding $Y(\mathfrak{gl}_{a})\otimes Y(\mathfrak{gl}_{n-a})\hookrightarrow Y(\mathfrak{gl}_{n})$
The off-shell Bethe vectors $\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})$ factorizes into the product of $\mathfrak{gl}_{a}$-type pre-Bethe vectors $\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})$ and $\mathfrak{gl}_{n-a}$-type pre-Bethe vectors $\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})$ acting on the vacuum vector $|0\rangle$:
$\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})=\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\,|0\rangle\,.$ (3.13)
@@
Let us consider the recurrence relations (3.10) for two different ranges of the indices $\ell$ and $k$: $1\leq\ell<k\leq a$ and $a+1\leq\ell<k\leq n$. These ranges for $k$ and $\ell$ ensure that the set ${\bar{t}}^{a}$ remains empty. Since $\mathbb{B}_{\mathfrak{gl}}(\varnothing)=|0\rangle$ (i.e. $\mathcal{B}_{\mathfrak{gl}}(\varnothing)=1$) the recurrence relations will prove the statement of (3.13) by induction on the cardinalities.
@@
Splitting property of the Bethe vectors. Let us define the operation $\mathcal{Z}^{a+1}_{a}(t^{a})$ as the operation which adds the parameter $t^{a}$ to the set ${\bar{t}}^{a}$ of the Bethe vector $\mathbb{B}_{\mathfrak{gl}}({\bar{t}})$ according to definition (3.4). A direct consequence of the proposition 3.2 and the rectangular recurrence relations (3.10) is the presentation of the Bethe vector $\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}^{a-1}_{1},t^{a},\{{\bar{t}}^{s}\}_{a+1}^{n-1})$ in the form
$\begin{split}&\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}^{a-1}_{1},t^{a},\{{\bar{t}}^{s}\}_{a+1}^{n-1})=\mathcal{Z}^{a+1}_{a}(t^{a})\Big{(}\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\Big{)}\,|0\rangle=\\ &\quad=\frac{1}{\mu^{k}_{\ell}(t^{a};{\bar{t}})}\ \sum_{i=1}^{a}\sum_{j=a+1}^{n}\sum_{\rm part}\Xi^{\ell,k}_{i,j}(t^{a};{\bar{t}})\ T_{i,j}(t^{a})\cdot\Big{(}\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\Big{)}\,|0\rangle\,.\end{split}$ (3.14)
@@ 4 Bethe vectors in $\mathfrak{o}_{2n+1}$-invariant models
The commutation relations between the zero mode operators ${\sf T}_{\ell+1,\ell}$ and the monodromy matrix elements take the form for $10\leq\ell\leq n-$ 1
@@
$\gamma_{\ell}(t_{1},t_{2})=\begin{cases}\displaystyle\mathfrak{f}(t_{1},t_{2})=\frac{t_{1}-t_{2}+c/2}{t_{1}-t_{2}},&\ell=0,\\[8.61108pt] \displaystyle\gamma(t_{1},t_{2})=\frac{g(t_{1},t_{2})}{h(t_{2},t_{1})}=\frac{c^{2}}{(t_{1}-t_{2})(t_{2}-t_{1}+c)},\quad&\ell=1,\ldots,n-1.\end{cases}$ (4.10)
@@ 4.1 Rectangular recurrence relations for $\mathfrak{o}_{2n+1}$
    $\sigma_{m}=2\Theta(m-1)-1=\begin{cases}1,\quad m>0,\\ -1,\quad m\leq 0\end{cases}$ (4.11)
@@
    $\begin{split}\Xi^{\ell,k}_{i,j}(z;{\bar{t}})&=\psi_{\ell}(z,{\bar{t}}_{\scriptscriptstyle\rm I})\ \phi_{k}(z,{\bar{t}}_{\scriptscriptstyle\rm III})\ \Gamma^{R}_{0,n}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{{\scriptscriptstyle\rm II},{\scriptscriptstyle\rm III}})\ \Gamma^{L}_{0,n}({\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{{\scriptscriptstyle\rm III}})=\\ &=\psi_{\ell}(z,{\bar{t}}_{\scriptscriptstyle\rm I})\ \phi_{k}(z,{\bar{t}}_{\scriptscriptstyle\rm III})\ \Gamma^{R}_{0,n}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{{\scriptscriptstyle\rm II}})\ \Gamma^{L}_{0,n}({\bar{t}}_{{\scriptscriptstyle\rm I},{\scriptscriptstyle\rm II}},{\bar{t}}_{{\scriptscriptstyle\rm III}})\,,\end{split}$ (4.20)
    where the equality between first and second lines in (4.20) directly follows from the definitions (4.9). Note that $\Gamma^{L}_{i,j}({\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{\scriptscriptstyle\rm I})=\Gamma^{R}_{i,j}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{\scriptscriptstyle\rm II})\equiv 1$ for $i\geq j$.
@@
For any $\ell,k$ such that $-n\leq l<k\leq n$ the off-shell Bethe vector $\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})$ satisfies the rectangular recurrence relation
@@
where the sum is over partitions ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}\}$ with cardinalities depending on the indices $-n\leq i\leq\ell<k\leq j\leq n$ and given by the equalities (4.16), (4.17) and (4.18). If the cardinality of the set ${\bar{t}}^{s}$ is small and according to (4.16) and (4.17), one may have $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|+|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|>|{\bar{t}}^{s}|$ for some $i$ and $j$. Then the terms in the r.h.s. of (4.21) for such values of the indices $i$ and $j$ have to be discarded.
Let us stress that, in opposition to the $\mathfrak{gl}_{n}$ case, for $\mathfrak{o}_{2n+1}$-invariant models, depending on the color $s$, the subsets ${\bar{t}}^{s}_{\scriptscriptstyle\rm I}$ and ${\bar{t}}^{s}_{\scriptscriptstyle\rm III}$ can be both non-empty at the same time, and the cardinalities of ${\bar{t}}^{s}_{\scriptscriptstyle\rm I}$ and ${\bar{t}}^{s}_{\scriptscriptstyle\rm III}$ can be up to 2.
@@ 4.2 Examples of recurrence relations
    $\begin{split}&\mathcal{Z}^{\ell+1}_{\ell}\cdot{\mathbb{B}}({\bar{t}})={\mathbb{B}}({\bar{t}}^{0},\ldots,{\bar{t}}^{\ell-1},\{{\bar{t}}^{\ell},z\},{\bar{t}}^{\ell+1},\ldots,{\bar{t}}^{n-1})=\\ &\quad=\sum_{i=-n}^{\ell}\sum_{j=\ell+1}^{n}\sum_{\rm part}\frac{g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\,g({\bar{t}}^{\ell+1}_{\scriptscriptstyle\rm III},z)}{\lambda_{\ell+1}(z)\,g(z,{\bar{t}}^{\ell-1})\,h({\bar{t}}^{\ell},z)\,h(z,{\bar{t}}^{\ell})\,g({\bar{t}}^{\ell+1},z)}\times\\ &\qquad\times\Gamma^{R}_{[i],\ell_{i}}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{{\scriptscriptstyle\rm II}})\ \Gamma^{L}_{\ell+1,j}({\bar{t}}_{{\scriptscriptstyle\rm I},{\scriptscriptstyle\rm II}},{\bar{t}}_{{\scriptscriptstyle\rm III}})\ \sigma_{i+1}\ T_{i,j}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})\,,\end{split}$ (4.22)
@@
    $\begin{split}{\mathbb{B}}(\{t^{s}\}_{0}^{n-2},\{{\bar{t}}^{n-1},z\})&=\frac{1}{\lambda_{n}(z)\ h(z,{\bar{t}}^{n-1})}\times\\ &\times\sum_{i=-n}^{n-1}\sum_{\rm part}\frac{\sigma_{i+1}\,T_{i,n}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})}{g(z,{\bar{t}}^{n-2}_{\scriptscriptstyle\rm II})\,h({\bar{t}}^{n-1}_{\scriptscriptstyle\rm II},z)}\prod_{s=0}^{n-1}\gamma_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II})\prod_{s=1}^{n-1}\frac{h({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II})}\end{split}$
@@
  • When $0<\ell<n$, we get a so called shifted recurrence relation for the off-shell Bethe vectors. It corresponds to an extension of the set of Bethe parameters ${\bar{t}}^{\ell}$ by the shifted parameter $z_{\ell}=z-c(\ell-1/2)$:
    $\begin{split}&\mathcal{Z}^{-\ell}_{-\ell-1}\cdot{\mathbb{B}}({\bar{t}})={\mathbb{B}}({\bar{t}}^{0},\ldots,{\bar{t}}^{\ell-1},\{{\bar{t}}^{\ell},z_{\ell}\},{\bar{t}}^{\ell+1},\ldots,{\bar{t}}^{n-1})=\\ &\quad=\sum_{i=-n}^{-\ell-1}\sum_{j=-\ell}^{n}\sum_{\rm part}\frac{(-1)^{\delta_{j\geq\ell+1}}\ g({\bar{t}}^{\ell+1}_{\scriptscriptstyle\rm I},z_{\ell+1})\,h(z_{\ell},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm III})\,g(z_{\ell-1},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm III})}{\lambda_{-\ell}(z)\,g({\bar{t}}^{\ell+1},z_{\ell+1})\,h({\bar{t}}^{\ell},z_{\ell})\,h(z_{\ell},{\bar{t}}^{\ell})\,g(z_{\ell-1},{\bar{t}}^{\ell-1})}\times\\ &\qquad\times\Gamma^{R}_{\ell+1,-i}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{{\scriptscriptstyle\rm II},{\scriptscriptstyle\rm III}})\ \Gamma^{L}_{[-j],\ell_{j}}({\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{{\scriptscriptstyle\rm III}})\ \sigma_{j}\ T_{i,j}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})\,,\end{split}$ (4.24)
    where $\ell_{j}={\rm max}(\ell,|j|)$, $[j]=\frac{1}{2}(j+|j|)$ as above, and the cardinalities of the subsets in the sum over partitions are
    $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=\begin{cases}0,&s<\ell+1,\\ \Theta(-i-s-1),&s\geq\ell+1,\end{cases}\qquad|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=\begin{cases}\Theta(s+j)+\Theta(j-s-1),&s<\ell,\\ \Theta(j-s-1),&s\geq\ell.\end{cases}$ (4.25)
    The sign factor $(-1)^{\delta_{j\geq\ell+1}}$ in the second line of (4.24) is due to the identity $g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm III},z_{\ell-1})^{-1}=-h(z_{\ell},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm III})$, valid when the cardinality of the subset ${\bar{t}}^{\ell}_{\scriptscriptstyle\rm III}$ is equal to one. According to (4.25) this happens when $j\geq\ell+1$.
    For $\ell=n-1$ the recurrence relation (4.24) can be rewritten in the form
@@
    and coincides exactly with the recurrence relation (3.19) from the paper [18]. The sign factor $(-1)^{\delta_{j\geq\ell+1}}$ turns in this case to the sign factor $(-1)^{\delta_{j,n}}$.
@@
For the Bethe vector $\mathcal{Z}^{1}_{-1}\cdot{\mathbb{B}}({\bar{t}}^{0})$, there is no summation over partitions and according to (4.15) it is equal to
$\mathcal{Z}^{1}_{-1}\cdot{\mathbb{B}}({\bar{t}}^{0})={\mathbb{B}}(\{{\bar{t}}^{0},z,z_{0}\})=-\frac{\lambda_{1}(z)}{2}\ T_{-1,1}(z)\cdot{\mathbb{B}}({\bar{t}}^{0})\,.$ (4.31)
@@
Remark 4.1.
The $n=1$ example shows that for $\mathfrak{o}_{2n+1}$-invariant integrable models, there is no simple symmetry transformation which relates the shifted (4.24) and non-shifted (4.22) recurrence relations. This is in opposition to the $\mathfrak{gl}_{n}$-integrable models, where there is a symmetry relation for Bethe vectors, see [21] and the $\mathfrak{gl}_{n}$ reduction in section 4.3.
Indeed, one can replace $z\to z_{0}$ in (4.32) so the l.h.s. of (4.32) and (4.33) coincide. But then, in order to prove that the r.h.s. of these equalities also coincide, one has to use the following relation between monodromy matrix entries
$\begin{split}T_{-1,0}(z)&=-T_{0,1}(z_{0})\,T_{1,1}(z_{0})^{-1}\,T_{0,0}(z)+\\ &\quad+\ \Big{(}T_{0,1}(z_{0})\,T_{1,1}(z_{0})^{-1}\,T_{0,1}(z)\,T_{1,1}(z)^{-1}+T_{-1,1}(z)\,T_{1,1}(z)^{-1}\Big{)}\,T_{1,0}(z)\,,\end{split}$ (4.34)
which is a direct consequence of (2.9). Then, in principle, using the action formulas for $T_{0,0}(z)$, $T_{1,1}(z)^{-1}$ and $T_{1,0}(z)$ one could connect the shifted and non-shifted recurrence formula, but these calculations are too complicated to be carried in the general case. Action formulas for $T_{0,0}(z)$, $T_{1,1}(z)$ and $T_{1,0}(z)$ have been computed in [21] for $n=1$ and in [18] for arbitrary $n$.
In conclusion, in $\mathfrak{o}_{2n+1}$-invariant quantum integrable models, there is no simple answer to the question of the existence of a symmetry transformation in the space of the off-shell Bethe vectors based on the symmetry relation (2.9) which relates the non-shifted and shifted recurrence relations.

### v2

@@ 3.3 Embedding $Y(\mathfrak{gl}_{a})\otimes Y(\mathfrak{gl}_{n-a})\hookrightarrow Y(\mathfrak{gl}_{n})$
The off-shell Bethe vectors $\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})$ has a representation as the product of $\mathfrak{gl}_{a}$ pre-Bethe vectors $\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})\in Y(\mathfrak{gl}_{a})$ and $\mathfrak{gl}_{n-a}$ pre-Bethe vectors $\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\in Y(\mathfrak{gl}_{n-a})$ acting on the vacuum vector $|0\rangle$:
$\begin{split}\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})&=\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\,|0\rangle\\ &=\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})\,|0\rangle\,.\end{split}$ (3.13)
@@
We will prove only the first equality in (3.13). The second equality can be proved analogously. Let us consider the recurrence relations (3.10) for two different ranges of the indices $\ell$ and $k$: $1\leq\ell<k\leq a$ and $a+1\leq\ell<k\leq n$. These ranges for $k$ and $\ell$ ensure that the set ${\bar{t}}^{a}$ remains empty. Since $\mathbb{B}_{\mathfrak{gl}}(\varnothing)=|0\rangle$ (i.e. $\mathcal{B}_{\mathfrak{gl}}(\varnothing)=1$) the recurrence relations will prove the statement of (3.13) by induction on the cardinalities.
@@
Splitting property of the Bethe vectors. Let us define the operation $\mathcal{Z}^{a+1}_{a}(t^{a}_{1})$ as the operation which adds the parameter $t^{a}_{1}$ of the color $a$ to the set ${\bar{t}}=\{\{{\bar{t}}^{s}\}^{a-1}_{1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1}\}$ of the Bethe vector $\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}^{a-1}_{1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})$ according to definition (3.4). A direct consequence of the proposition 3.2 and the rectangular recurrence relations (3.10) is the presentation of the Bethe vector $\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}^{a-1}_{1},t^{a}_{1},\{{\bar{t}}^{s}\}_{a+1}^{n-1})$ in the form
$\begin{split}&\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}^{a-1}_{1},\{t^{a}_{1}\},\{{\bar{t}}^{s}\}_{a+1}^{n-1})=\mathcal{Z}^{a+1}_{a}(t^{a}_{1})\Big{(}\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\Big{)}\,|0\rangle=\\ &\quad=\frac{1}{\mu^{a+1}_{a}(t^{a}_{1};{\bar{t}})}\ \sum_{i=1}^{a}\sum_{j=a+1}^{n}\sum_{\rm part}\Xi^{\ell,k}_{i,j}(t^{a}_{1};{\bar{t}})\ T_{i,j}(t^{a}_{1})\cdot\Big{(}\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\Big{)}\,|0\rangle\,.\end{split}$ (3.14)
@@ 4 Bethe vectors in $\mathfrak{o}_{2n+1}$-invariant models
The commutation relations (2.12) between the zero mode operators ${\sf T}_{\ell+1,\ell}$ and the monodromy matrix elements take the form for $10\leq\ell\leq n-$ 1
@@
$\gamma_{\ell}(u,v)=\begin{cases}\displaystyle\mathfrak{f}(u,v)=\frac{u-v+c/2}{u-v},&\ell=0,\\[8.61108pt] \displaystyle\gamma(u,v)=\frac{g(u,v)}{h(v,u)}=\frac{c^{2}}{(u-v)(v-u+c)},\quad&\ell=1,\ldots,n-1.\end{cases}$ (4.10)
@@ 4.1 Rectangular recurrence relations for $\mathfrak{o}_{2n+1}$
    $\sigma_{m}=2\Theta(m-1)-1=\begin{cases}1,\quad&m>0,\\ -1,\quad&m\leq 0\end{cases}$ (4.11)
@@
    $\begin{split}\Xi^{\ell,k}_{i,j}(z;{\bar{t}})&=\psi_{\ell}(z,{\bar{t}}_{\scriptscriptstyle\rm I})\ \phi_{k}(z,{\bar{t}}_{\scriptscriptstyle\rm III})\ \Gamma^{R}_{0,n}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{{\scriptscriptstyle\rm II},{\scriptscriptstyle\rm III}})\ \Gamma^{L}_{0,n}({\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{{\scriptscriptstyle\rm III}})\\ &=\psi_{\ell}(z,{\bar{t}}_{\scriptscriptstyle\rm I})\ \phi_{k}(z,{\bar{t}}_{\scriptscriptstyle\rm III})\ \Gamma^{R}_{0,n}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{{\scriptscriptstyle\rm II}})\ \Gamma^{L}_{0,n}({\bar{t}}_{{\scriptscriptstyle\rm I},{\scriptscriptstyle\rm II}},{\bar{t}}_{{\scriptscriptstyle\rm III}})\,,\end{split}$ (4.20)
    where the equality between the first and second lines in (4.20) directly follows from the definitions (4.9), with the notation ${\bar{t}}_{\scriptscriptstyle\rm I,\scriptscriptstyle\rm II}=\{{\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{\scriptscriptstyle\rm II}\}$, and ${\bar{t}}_{\scriptscriptstyle\rm II,\scriptscriptstyle\rm III}=\{{\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{\scriptscriptstyle\rm III}\}$. Note that $\Gamma^{L}_{i,j}({\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{\scriptscriptstyle\rm I})=\Gamma^{R}_{i,j}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{\scriptscriptstyle\rm II})\equiv 1$ for $i\geq j$.
@@
For any $\ell,k$ such that $-n\leq\ell<k\leq n$ the off-shell Bethe vector $\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})$ satisfies the rectangular recurrence relation
@@
where the sum is over partitions ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}\}$ with cardinalities depending on the indices $-n\leq i\leq\ell<k\leq j\leq n$ and given by the equalities (4.16), (4.17) and (4.18).
If the cardinality of the set ${\bar{t}}^{s}$ is small and according to (4.16) and (4.17), one may have $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|+|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|>|{\bar{t}}^{s}|$ for some $i$ and $j$. Then the terms in the r.h.s. of (4.21) for such values of the indices $i$ and $j$ have to be discarded.
Let us stress that, in opposition to the $\mathfrak{gl}_{n}$ case, for $\mathfrak{o}_{2n+1}$-invariant models, depending on the color $s$, the subsets ${\bar{t}}^{s}_{\scriptscriptstyle\rm I}$ and ${\bar{t}}^{s}_{\scriptscriptstyle\rm III}$ satisfy $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|+|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|\leq 2$.
@@ 4.2 Examples of recurrence relations
    $\begin{split}&\mathcal{Z}^{\ell+1}_{\ell}\cdot{\mathbb{B}}({\bar{t}})={\mathbb{B}}({\bar{t}}^{0},\ldots,{\bar{t}}^{\ell-1},\{{\bar{t}}^{\ell},z\},{\bar{t}}^{\ell+1},\ldots,{\bar{t}}^{n-1})=\\ &\quad=\sum_{i=-n}^{\ell}\sum_{j=\ell+1}^{n}\sum_{\rm part}\frac{g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\,g({\bar{t}}^{\ell+1}_{\scriptscriptstyle\rm III},z)}{\lambda_{\ell+1}(z)\,g(z,{\bar{t}}^{\ell-1})\,h({\bar{t}}^{\ell},z)\,h(z,{\bar{t}}^{\ell})\,g({\bar{t}}^{\ell+1},z)}\times\\ &\qquad\qquad\qquad\times\Gamma^{R}_{[i],\ell_{i}}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{{\scriptscriptstyle\rm II}})\ \Gamma^{L}_{\ell+1,j}({\bar{t}}_{{\scriptscriptstyle\rm I},{\scriptscriptstyle\rm II}},{\bar{t}}_{{\scriptscriptstyle\rm III}})\ \sigma_{i+1}\ T_{i,j}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})\,,\end{split}$ (4.22)
@@
    $\begin{split}{\mathbb{B}}(\{{\bar{t}}^{s}\}_{0}^{n-2},\{{\bar{t}}^{n-1},z\})&=\frac{1}{\lambda_{n}(z)\ h(z,{\bar{t}}^{n-1})}\times\\ &\times\sum_{i=-n}^{n-1}\sum_{\rm part}\frac{\sigma_{i+1}\,T_{i,n}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})}{g(z,{\bar{t}}^{n-2}_{\scriptscriptstyle\rm II})\,h({\bar{t}}^{n-1}_{\scriptscriptstyle\rm II},z)}\prod_{s=0}^{n-1}\gamma_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II})\prod_{s=1}^{n-1}\frac{h({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II})}\end{split}$
@@
  • The cases of $-n\leq\ell<-1$ corresponds to so called shifted recurrence relation for the off-shell Bethe vectors. To describe these cases we set $\ell=-l-1$ with $0<l<n$. They correspond to an extension of the set of Bethe parameters ${\bar{t}}^{l}$ by the shifted parameter $z_{l}=z-c(l-1/2)$:
    $\begin{split}&\mathcal{Z}^{-l}_{-l-1}\cdot{\mathbb{B}}({\bar{t}})={\mathbb{B}}({\bar{t}}^{0},\ldots,{\bar{t}}^{l-1},\{{\bar{t}}^{l},z_{l}\},{\bar{t}}^{l+1},\ldots,{\bar{t}}^{n-1})=\\ &\quad=\sum_{i=-n}^{-l-1}\sum_{j=-l}^{n}\sum_{\rm part}\frac{(-1)^{\delta_{j\geq l+1}}\ g({\bar{t}}^{l+1}_{\scriptscriptstyle\rm I},z_{l+1})\,h(z_{l},{\bar{t}}^{l}_{\scriptscriptstyle\rm III})\,g(z_{l-1},{\bar{t}}^{l-1}_{\scriptscriptstyle\rm III})}{\lambda_{-l}(z)\,g({\bar{t}}^{l+1},z_{l+1})\,h({\bar{t}}^{l},z_{l})\,h(z_{l},{\bar{t}}^{l})\,g(z_{l-1},{\bar{t}}^{l-1})}\times\\ &\qquad\qquad\qquad\times\Gamma^{R}_{l+1,-i}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{{\scriptscriptstyle\rm II},{\scriptscriptstyle\rm III}})\ \Gamma^{L}_{[-j],l_{j}}({\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{{\scriptscriptstyle\rm III}})\ \sigma_{j}\ T_{i,j}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})\,,\end{split}$ (4.24)
    where $l_{j}={\rm max}(l,|j|)$, $[j]=\frac{1}{2}(j+|j|)$ as above, and the cardinalities of the subsets in the sum over partitions are
    $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=\begin{cases}0,&s<l+1,\\ \Theta(-i-s-1),&s\geq l+1,\end{cases}\qquad|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=\begin{cases}\Theta(s+j)+\Theta(j-s-1),&s<l,\\ \Theta(j-s-1),&s\geq l.\end{cases}$ (4.25)
    The sign factor $(-1)^{\delta_{j\geq l+1}}$ in the second line of (4.24) is due to the identity $g({\bar{t}}^{l}_{\scriptscriptstyle\rm III},z_{l-1})^{-1}=-h(z_{l},{\bar{t}}^{l}_{\scriptscriptstyle\rm III})$, valid when the cardinality of the subset ${\bar{t}}^{l}_{\scriptscriptstyle\rm III}$ is equal to one. According to (4.25) this happens when $j\geq l+1$.
    For $l=n-1$ the recurrence relation (4.24) can be rewritten in the form
@@
    and coincides exactly with the recurrence relation (3.19) from the paper [19]. The sign factor $(-1)^{\delta_{j\geq l+1}}$ turns in this case to the sign factor $(-1)^{\delta_{j,n}}$.
@@
For the Bethe vector $\mathcal{Z}^{1}_{-1}\cdot{\mathbb{B}}({\bar{t}}^{0})$, there is no summation over partitions and according to (4.1) and (4.3) it is equal to
$\mathcal{Z}^{1}_{-1}\cdot{\mathbb{B}}({\bar{t}}^{0})={\mathbb{B}}(\{{\bar{t}}^{0},z,z_{0}\})=-\frac{2}{\lambda_{1}(z)}\ T_{-1,1}(z)\cdot{\mathbb{B}}({\bar{t}}^{0})\,.$ (4.31)